# Searching for leptophilic composite asymmetric dark sector at $e^{+}e^{-}$ colliders

### v1

@@ III Existing collider constraints on model parameters
Dark quark pairs can also be produced via $t$-channel mediator exchange at the LEP. Following dark showers and hadronization, if all dark mesons escape the detector, they contribute to missing energy. With initial state radiation (ISR), this also yields a mono-photon signature. For $10\text{ MeV}\leq m_{\pi_{d}}<2m_{\mu}$, we conservatively set $m_{\pi_{d}}=f_{\pi_{d}}=0.2\text{ GeV}$ and require $\pi_{d}$ proper decay length greater than $12$ m to ensure that all dark mesons escape the detector volume of LEP. The black dashed line in Fig. 3 gives upper bounds $\kappa\sim\mathcal{O}(1-10)$. We further recast mono-photon constraints at the LEP from the $t$-channel scalar mediator and Dirac fermionic DM model with the cut-off scale $\Lambda$ (Eq. 4) and the purple line in Fig. 4 ($\bar{\chi}e\bar{e}\chi$) of Ref. Fox:2011fx to the leptophilic composite ADM via:
@@
where the factor of 3 originates from $SU(3)_{D}$ charges. Assuming light dark quarks and determining lower bounds of $\Lambda$, the black regions in Fig. 3 indicate where predicted cross-sections exceed experimental bounds, consistent with the $1/\Lambda^{2}$ scaling.
Finally, a dark pion lifetime exceeding one second would violate observed light-element abundances and the CMB spectrum, as constrained by Big Bang Nucleosynthesis (BBN) and cosmic microwave background (CMB) measurements Lindley:1984bg ; Dev:2025pru . We therefore incorporate these constraints by excluding regions where the proper lifetime of $\pi_{d}$ is greater than 1 sec, indicated by the green shaded areas in Figs. 3 and 4. Notably, only a narrow region of parameter space remains in Fig. 3 for $m_{X}\lesssim 1$ TeV.
@@ IV Sparse dark showers at BaBar, Belle II and GAZELLE
In this and the following section, we employ FeynRules Alloul:2013bka to generate a UFO model file for Eq. (1), and then parton-level Monte Carlo events are generated using MadGraph5_aMC@NLO Alwall:2014hca . Subsequently, the Hidden Valley module within Pythia8.313 Sjostrand:2014zea simulates dark parton showering and hadronization processes. Parameter settings closely follow Refs. Schwaller:2015gea ; Mies:2020mzw ; Archer-Smith:2021ntx ; Carrasco:2023loy , specifically: $N_{C_{d}}=3$, $N_{f_{d}}=7$, $\Lambda_{d}=m_{q_{d}}=2m_{\pi_{d}}$, and $m_{\rho_{d}}=4m_{\pi_{d}}$. In our simulation, diagonal and off-diagonal HV-mesons share identical decay patterns. We implement the running coupling $\alpha_{d}$ and apply a $p_{T}$ cutoff of $1.1\Lambda_{d}$ to dark showers. The fraction of vector and pseudoscalar HV-mesons is calculated by following Appendix A of Ref. Knapen:2021eip . Finally, the ratio of dark pion mass over dark pion decay constant, $m_{\pi_{d}}/f_{\pi_{d}}$, can range from $\mathcal{O}(1-10)$ Berlin:2018tvf ; Choi:2018iit ; Kuwahara:2023vfc ; Bernreuther:2023kcg . For this analysis, we fix $f_{\pi_{d}}=m_{\pi_{d}}$. This choice only affects the dark pion decay width, which can be easily rescaled using Eq. (3).
@@ IV.1 Recasting LLP searches at the BaBar
The detector resolution is performed by the smearing effect of the relevant uncertainties of observables. Since recasting efficiencies are only available for LLPs with $m>0.5~\text{GeV}$ and $c\tau>5~\text{mm}$ as noted in Refs. BaBar:2015jvu ; Bernreuther:2022jlj , we restrict analysis to two of our BPs: $m_{\pi_{d}}=0.7~\text{GeV}$ and $1~\text{GeV}$, requiring $c\tau>5~\text{mm}$ for dark pions.
@@
Figure 5: Kinematic distributions for the signal benchmark points BP2 and BP3 at the BaBar. The distributions of the leading $p_{T}$ and subleading $p_{T}$ dark pions and their daughter muons are represented by solid and dashed lines, respectively. Top-left: transverse momentum $p_{T}$ of muons from dark pion decays; Top-right: invariant mass $m_{\mu^{+}\mu^{-}}$ of muon pairs; Bottom-left: transverse decay distance $R_{xy}$ of dark pions; Bottom-right: track impact parameter significance $d_{0}/\sigma_{0}$.
@@
Figure 5 shows the relevant kinematic distributions of these two signal BPs. First, the transverse momentum distribution of the muon, $P_{T}(\mu)$ (top-left), is generally larger for heavier $\pi_{d}$ masses. This is because, under a lower total $\pi_{d}$ yield, each $\pi_{d}$ carries higher average energy, leading to muons from its decay with larger $p_{T}$. Second, in the $\mu^{+}\mu^{-}$ invariant mass distribution, $m_{\mu^{+}\mu^{-}}$ (top-right), the position of the reconstructed mass peak directly reflects the mass of $\pi_{d}$. Regarding the transverse decay length, $R_{xy}$ (bottom-left), since $\pi_{d}$ is long-lived, its decay vertex is significantly displaced from the IP in the transverse plane. Moreover, due to the proper decay length $c\tau_{0}\propto 1/m_{\pi_{d}}^{3}$, a lighter $\pi_{d}$ has a longer decay length, resulting in an $R_{xy}$ distribution that extends to larger values and exhibits a more pronounced long tail for BP2 compared to BP3. Finally, the transverse impact parameter $d_{0}$ (bottom-right) further characterizes the long-lived nature of $\pi_{d}$. The value of $d_{0}$ is significantly larger than the uncertainty $\sigma_{0}$ which clearly distinguishes $\pi_{d}$ decays from prompt decays.
@@
Figure 6: Exclusion limits for three benchmark points, $m_{\pi_{d}}=0.32$, $0.7$ and $1$ GeV for three experiments: BaBar, Belle II and L-GAZELLE.
@@
The production cross-section for $e^{+}e^{-}\to\overline{q_{d}}q_{d}$ scales as $(\kappa/m_{X})^{4}$, while the proper lifetime of $\pi_{d}$ is inversely proportional to this factor. We therefore recast BaBar LLP search constraints onto the $(m_{X},\kappa)$ parameter plane. Specifically, we extract exclusion limits from the $\mu^{+}\mu^{-}$ channel (red curve in Fig. 2 of Ref. BaBar:2015jvu ), plotting LLP mass versus $\sigma(e^{+}e^{-}\to LX)\mathcal{B}(L\to f)\epsilon(f)$, where $\sigma(e^{+}e^{-}\to LX)$ is the LLP production cross-section ($L$ is the LLP candidate, and $X$ are associated particles); $\mathcal{B}(L\to f)$ is the branching fraction to final state of a pair of $f$ particles (a muon pair in our case); $\epsilon(f)$ is the detection efficiency to such final state. For $m_{\pi_{d}}=0.7$ GeV and $1$ GeV, exclusion limits appear in gray regions of Fig. 6. The upper boundaries correspond to short dark pion lifetimes that preclude LLP classification, while lower boundaries indicate long-lived pions escaping the detector’s sensitive volume. The inverse relationship of constraints between $m_{X}$ and $\kappa$ is clearly evident in Fig. 6.
@@ IV.2 Detecting dark showers at the Belle II
After Cuts 13.7413.74% 65.0165.01% 51.6251.62%
Displaced Vertex 93.5693.56% 79.4179.41% 46.2446.24%
After Cuts 15.5615.56% 60.1260.12% 35.1635.16%
@@
Figure 7: Distributions for leading (solid line) and subleading (dashed line) $p_{T}$ dark pions ($\pi_{d}$) in BP1, BP2, and BP3 at the Belle II. Top-left: Number of dark pions per event, $N_{\pi_{d}}$; Top-right: Transverse decay length, $R_{xy}$, of leading and subleading $\pi_{d}$’s; Bottom-left: Longitudinal flight distance, $R_{z}$; Bottom-right: Angle between the decay vertex vector and beam axis, $\theta_{\text{lab}}$.
@@
Figure 8: The kinematic distributions for muons from leading $p_{T}$ (solid line) and subleading $p_{T}$ (dashed line) dark pions in BP1, BP2, and BP3 at the Belle II. Top-left: Transverse momentum of muons, $p_{T}(\mu)$; Top-right: Invariant mass of muon pairs, $m_{\mu^{+}\mu^{-}}$; Bottom-left: Energy of muon pairs, $E_{\mu^{+}\mu^{-}}$; Bottom-right: Angular separation between opposite-sign muons, $\Delta R_{\mu^{+}\mu^{-}}$.
@@
We further apply the event selection criteria summarized in Table 3 to identify the signal events and suppress the background ones. An event is selected if at least two leading-$p_{T}$ $\pi_{d}$ particles satisfy the following conditions: $0.2~\text{cm}<R_{xy}<60~\text{cm}$, $55~\text{cm}\leq z\leq 140~\text{cm}$, and $17^{\circ}\leq\theta_{\text{lab}}\leq 150^{\circ}$. Additionally, their daughter muons must satisfy $p_{T}>0.05~\text{GeV}$, and the reconstructed invariant mass of the muon pair must fulfill $m_{\mu^{+}\mu^{-}}<0.48~\text{GeV}$ or $m_{\mu^{+}\mu^{-}}>0.52~\text{GeV}$.
@@
Table 4 summarizes the efficiencies of three trigger types and after applying event selection criteria for three BPs. Overall, the efficiencies of all three triggers exhibit significant variations with $m_{\pi_{d}}$, and the relative ordering of efficiencies among the BPs changes considerably after the same selection is applied. Specifically, the efficiency of Two tracks trigger decreases as $m_{\pi_{d}}$ increases, dropping from $75.82\%$ for BP1 to $34.16\%$ for BP3. This behavior can be understood from the distribution of $\theta_{\text{lab}}$. For $m_{\pi_{d}}=1~\text{GeV}$, the tracks from the leading and sub-leading $p_{T}$ $\pi_{d}$’s are more dispersed in $\theta_{\text{lab}}$ distribution, with many falling outside the trigger acceptance region of $38^{\circ}\leq\theta_{\text{lab}}\leq 127^{\circ}$. In this case, the two selected tracks are very likely to originate from the same dark pion, and their opening angle $\Delta\phi$ rarely exceeds $\pi/2$, leading to the lowest efficiency. In contrast, for $m_{\pi_{d}}=0.32$ and $0.7~\text{GeV}$, the $\theta_{\text{lab}}$ distribution is more concentrated, and the higher $\pi_{d}$ multiplicity results in higher efficiency. One muon trigger maintains high efficiency across all three BPs, peaking at $89.22\%$ for BP2. The efficiency of Displaced vertex trigger decreases with increasing $m_{\pi_{d}}$. This condition strongly depends on the decay length of $\pi_{d}$. The proper decay lengths for the three BPs are $c\tau_{0}=87.87~\text{mm}$ (BP1), $8.40~\text{mm}$ (BP2), and $2.88~\text{mm}$ (BP3), respectively. The proper decay length of BP1 is an order of magnitude larger than that of BP2, after multiplying by $\beta\gamma$ in the lab frame, it often exceeds the upper limit of the trigger condition. However, since the proper decay length follows an exponential distribution in the Monte Carlo simulation, the high multiplicity of BP1 enhances the probability that at least one displaced vertex satisfies the trigger condition. In contrast, for $m_{\pi_{d}}=1~\text{GeV}$, the $R_{xy}$ distribution of the two leading $p_{T}$ $\pi_{d}$’s is concentrated in a very small region, resulting in the fewest events satisfying the trigger condition. After applying the cuts in Table 3, the order of efficiencies in these three cases may change. From the $R_{xy}$ distribution of $\pi_{d}$’s, for $m_{\pi_{d}}=0.32~\text{GeV}$, many events have decay lengths for both $\pi_{d}$’s exceeding 60 cm, leading to a significant reduction in efficiency. In comparison, for $m_{\pi_{d}}=0.7~\text{GeV}$, most events fall within the required range, resulting in the highest efficiency.
@@
We then analyze relevant signal kinematic distributions for $\pi_{d}$’s and muons at the Belle II in Figs. 7 and 8. First, as the $m_{\pi_{d}}$ increases from BP1 to BP3, the $N_{\pi_{d}}$ distribution (top-left) decreases from approximately 6 to 2. Regarding the decay length, $R_{xy}$ and $R_{z}$ represent the transverse and longitudinal decay lengths of the dark pions in the lab frame, respectively. According to Eq. (4), under fixed parameters and with $f_{\pi_{d}}=m_{\pi_{d}}$, the proper decay length satisfies $c\tau_{0}\propto 1/m_{\pi_{d}}^{3}$. Therefore, lighter $\pi_{d}$’s have longer decay lengths, and their $R_{xy}$ (top-right) and $R_{z}$ (bottom-left) distributions exhibit larger values and longer tails. Finally, the $\theta_{\rm lab}(\text{track})$ distribution (bottom-right) is significantly dispersed for the BP1, while the other two BPs exhibit relatively concentrated distributions. By comparing $E$ and $p_{T}$ of daughter particles, it can be seen that for BP1 and BP2, the leading and subleading $p_{T}$ $\pi_{d}$’s have similar $p_{z}$ values, leading to small differences in their $\Delta\theta_{\rm lab}$ distributions. In contrast, for the BP3, although its $p_{T}$ distribution is similar, the leading $\pi_{d}$ carries higher energy, resulting in a larger $p_{z}$ difference. The greater difference in boost further leads to a significant separation in their $\theta_{\rm lab}$ distributions.
@@
As shown in Fig. 8, the $p_{T}(\mu)$ distribution (top-left) is similar to that in the BaBar analysis. In the $m_{\mu^{+}\mu^{-}}$ distribution (top-right), the reconstructed mass peaks clearly appear near the respective $m_{\pi_{d}}$: $0.32$ GeV, $0.7$ GeV, and $1$ GeV for BP1, BP2, and BP3, respectively, but exhibit different levels of broadening. The width of these distributions is influenced by momentum resolutions of muons Adachi:2018qme ; Kang:2021oes :
@@
the width of the $m_{\mu^{+}\mu^{-}}$ distribution increases from BP1 to BP3 as the momentum resolution grows with the $p_{T}(\mu)$. The $E_{\mu^{+}\mu^{-}}$ distribution (bottom-left) shows a clear $m_{\pi_{d}}$ and $p_{T}(\mu)$ dependence: muon pairs from the decay of heavier and higher-$p_{T}$ $\pi_{d}$’s carry more energy. Finally, in the $\Delta R_{\mu^{+}\mu^{-}}$ distribution (bottom-right), the average angular separation between the muons is smaller for higher $\pi_{d}$ multiplicity, while in BP3, which has fewer $\pi_{d}$’s, the $\Delta R$ distribution is generally larger.
@@
Since the coupling $\kappa$ and mediator mass $m_{X}$ affect both the $e^{+}e^{-}\to\overline{q_{d}}q_{d}$ production cross-section and $\pi_{d}$ lifetime, we map exclusion limits in the parameter space of $(m_{X},\kappa)$ plane. In particular, $\sigma(e^{+}e^{-}\to\overline{q_{d}}q_{d})$ and $\Gamma_{\pi_{d}}$ both scale as $(\kappa/m_{X})^{4}$. Applying the trigger conditions and event selection criteria (Tables 2 and 3), we obtain the signal efficiency $\epsilon_{S}$. Here an event is considered to pass the trigger selection if at least one of the three trigger conditions is satisfied. The expected signal yield is $N_{S}=\mathcal{L}\times\sigma(e^{+}e^{-}\to\overline{q_{d}}q_{d})\times\epsilon_{S}$, assuming optimistic integrated luminosity $\mathcal{L}=50~\text{ab}^{-1}$ for the Belle II experiment. With a background-free assumption, $N_{S}=2.3$ events establish the exclusion contours of the $90\%$ confidence level (CL) which correspond to the lower boundaries with purple solid lines in Fig. 6.
@@
As expected, small $\kappa/m_{X}$ yields a tiny cross-section and excessive $\pi_{d}$ decay length, reducing $N_{S}$ below detection ability of the Belle II experiment. Conversely, large $\kappa/m_{X}$ produces prompt $\pi_{d}$ decays without the signature of displaced vertices. In this situation, the efficiency drops sharply, necessitating enormous Monte Carlo samples. Therefore, we conservatively set $N_{S}=10$ to estimate upper boundaries with purple dashed lines in Fig. 6333Even with this relaxed requirement, more than $10^{8}$ simulated events are required for this estimation.. We can observe that the projected Belle II constraints also indicate inverse proportionality between $m_{X}$ and $\kappa$ for these three BPs. Notably, blank regions in the upper-left corner of $(m_{X},\kappa)$ plane could be covered by searching for multi-muons from the prompt decay of dark pions. However, detailed signal-background analysis is beyond the scope of this work.

### v2

@@ III Existing collider constraints on model parameters
Dark quark pairs can also be produced via $t$-channel mediator exchange at the LEP. Following dark showers and hadronization, if all dark mesons escape the detector, they contribute to missing energy. With initial-state radiation, this also yields a monophoton signature. For $10\text{ MeV}\leq m_{\pi_{d}}<2m_{\mu}$, we conservatively set $m_{\pi_{d}}=f_{\pi_{d}}=0.2\text{ GeV}$ and require $\pi_{d}$ proper decay length greater than $12$ m to ensure that all dark mesons escape the detector volume of LEP. The black dashed line in Fig. 3 gives upper bounds $\kappa\sim\mathcal{O}(1-10)$. We further recast monophoton constraints at the LEP from the $t$-channel scalar mediator and Dirac fermionic DM model with the cutoff scale $\Lambda$ [Eq. (4)] and the purple line in Fig. 4 ($\bar{\chi}e\bar{e}\chi$) of Ref. Fox:2011fx to the leptophilic composite ADM via
@@
where the factor of 3 originates from $SU(3)_{D}$ charges. Assuming light dark quarks and determining lower bounds of $\Lambda$, the black regions in Fig. 3 indicate where predicted cross sections exceed experimental bounds, consistent with the $1/\Lambda^{2}$ scaling.
Finally, a dark pion lifetime exceeding one second would violate observed light-element abundances and the cosmic microwave background (CMB) spectrum, as constrained by big bang nucleosynthesis and CMB measurements Lindley:1984bg ; Dev:2025pru . We, therefore, incorporate these constraints by excluding regions where the proper lifetime of $\pi_{d}$ is greater than one second, indicated by the green shaded areas in Figs. 3 and 4. Notably, only a narrow region of parameter space remains in Fig. 3 for $m_{X}\lesssim 1$ TeV.
@@ IV Sparse dark showers at BaBar, Belle II and GAZELLE
In this and the following section, we employ FeynRules Alloul:2013bka to generate a UFO model file for Eq. (1), and then parton-level Monte Carlo events are generated using MadGraph5_aMC@NLO Alwall:2014hca . Subsequently, the Hidden Valley (HV) module within Pythia8.313 Sjostrand:2014zea simulates dark parton showering and hadronization processes. Parameter settings closely follow Refs. Schwaller:2015gea ; Mies:2020mzw ; Archer-Smith:2021ntx ; Carrasco:2023loy , specifically, $N_{C_{d}}=3$, $N_{f_{d}}=7$, $\Lambda_{d}=m_{q_{d}}=2m_{\pi_{d}}$, and $m_{\rho_{d}}=4m_{\pi_{d}}$. In our simulation, diagonal and off-diagonal HV-mesons share identical decay patterns. We implement the running coupling $\alpha_{d}$ and apply a $p_{T}$ cutoff of $1.1\Lambda_{d}$ to dark showers. The fraction of vector and pseudoscalar HV-mesons is calculated by following Appendix A of Ref. Knapen:2021eip . Finally, the ratio of dark pion mass over dark pion decay constant, $m_{\pi_{d}}/f_{\pi_{d}}$, can range from $\mathcal{O}(1-10)$ Berlin:2018tvf ; Choi:2018iit ; Kuwahara:2023vfc ; Bernreuther:2023kcg . For this analysis, we fix $f_{\pi_{d}}=m_{\pi_{d}}$. This choice only affects the dark pion decay width, which can be easily rescaled using Eq. (3).
@@ IV.1 Recasting LLP searches at the BaBar
The detector resolution is performed by the smearing effect of the relevant uncertainties of observables. Since recasting efficiencies are only available for LLPs with $m>0.5~\text{GeV}$ and $c\tau>5~\text{mm}$, as noted in Refs. BaBar:2015jvu ; Bernreuther:2022jlj , we restrict analysis to two of our BPs: $m_{\pi_{d}}=0.7~\text{GeV}$ and $1~\text{GeV}$, requiring $c\tau>5~\text{mm}$ for dark pions.
@@
Figure 5: Kinematic distributions for the signal benchmark points BP2 and BP3 at the BaBar. The distributions of the leading $p_{T}$ and subleading $p_{T}$ dark pions and their daughter muons are represented by solid and dashed lines, respectively. Top left: transverse momentum $p_{T}$ of muons from dark pion decays. Top right: invariant mass $m_{\mu^{+}\mu^{-}}$ of muon pairs. Bottom left: transverse decay distance $R_{xy}$ of dark pions. Bottom-right: track impact parameter significance $d_{0}/\sigma_{0}$.
@@
Figure 5 shows the relevant kinematic distributions of these two signal BPs. First, the transverse momentum distribution of the muon, $P_{T}(\mu)$ (top left), is generally larger for heavier $\pi_{d}$ masses. This is because, under a lower total $\pi_{d}$ yield, each $\pi_{d}$ carries higher average energy, leading to muons from its decay with larger $p_{T}$. Second, in the $\mu^{+}\mu^{-}$ invariant mass distribution, $m_{\mu^{+}\mu^{-}}$ (top right), the position of the reconstructed mass peak directly reflects the mass of $\pi_{d}$. Regarding the transverse decay length, $R_{xy}$ (bottom left), since $\pi_{d}$ is long-lived, its decay vertex is significantly displaced from the IP in the transverse plane. Moreover, due to the proper decay length $c\tau_{0}\propto 1/m_{\pi_{d}}^{3}$, a lighter $\pi_{d}$ has a longer decay length, resulting in an $R_{xy}$ distribution that extends to larger values and exhibits a more pronounced long tail for BP2 compared to BP3. Finally, the transverse impact parameter $d_{0}$ (bottom right) further characterizes the long-lived nature of $\pi_{d}$. The value of $d_{0}$ is significantly larger than the uncertainty $\sigma_{0}$ which clearly distinguishes $\pi_{d}$ decays from prompt decays.
@@
Figure 6: Exclusion limits for three benchmark points, $m_{\pi_{d}}=0.32$, $0.7$ and $1$ GeV, for three experiments: BaBar, Belle II and L-GAZELLE.
@@
The production cross section for $e^{+}e^{-}\to\overline{q_{d}}q_{d}$ scales as $(\kappa/m_{X})^{4}$, while the proper lifetime of $\pi_{d}$ is inversely proportional to this factor. We, therefore, recast BaBar LLP search constraints onto the $(m_{X},\kappa)$ parameter plane. Specifically, we extract exclusion limits from the $\mu^{+}\mu^{-}$ channel (red curve in Fig. 2 of Ref. BaBar:2015jvu ), plotting LLP mass versus $\sigma(e^{+}e^{-}\to LX)\mathcal{B}(L\to f)\epsilon(f)$, where $\sigma(e^{+}e^{-}\to LX)$ is the LLP production cross section ($L$ is the LLP candidate, and $X$ are associated particles), $\mathcal{B}(L\to f)$ is the branching fraction to final state of a pair of $f$ particles (a muon pair in our case), and $\epsilon(f)$ is the detection efficiency to such final state. For $m_{\pi_{d}}=0.7$ GeV and $1$ GeV, exclusion limits appear in gray regions of Fig. 6. The upper boundaries correspond to short dark pion lifetimes that preclude LLP classification, while lower boundaries indicate long-lived pions escaping the detector’s sensitive volume. The inverse relationship of constraints between $m_{X}$ and $\kappa$ is clearly evident in Fig. 6.
@@ IV.2 Detecting dark showers at the Belle II
After Cuts 9.939.93% 65.0165.01% 51.6251.62%
Displaced Vertex 91.2491.24% 79.4179.41% 46.2446.24%
After Cuts 11.3911.39% 60.1260.12% 35.1635.16%
@@
Figure 7: Distributions for leading (solid line) and subleading (dashed line) $p_{T}$ dark pions ($\pi_{d}$) in BP1, BP2, and BP3 at the Belle II. Top left: number of dark pions per event, $N_{\pi_{d}}$. Top right: transverse decay length, $R_{xy}$, of leading and subleading $\pi_{d}$’s. Bottom left: longitudinal flight distance, $R_{z}$. Bottom right: angle between the decay vertex vector and beam axis, $\theta_{\text{lab}}$.
@@
Figure 8: The kinematic distributions for muons from leading $p_{T}$ (solid line) and subleading $p_{T}$ (dashed line) dark pions in BP1, BP2, and BP3 at the Belle II. Top left: transverse momentum of muons, $p_{T}(\mu)$. Top right: invariant mass of muon pairs, $m_{\mu^{+}\mu^{-}}$. Bottom left: energy of muon pairs, $E_{\mu^{+}\mu^{-}}$. Bottom-right: angular separation between opposite-sign muons, $\Delta R_{\mu^{+}\mu^{-}}$.
@@
We further apply the event selection criteria summarized in Table 3 to identify the signal events and suppress the background ones. An event is selected if at least two leading $p_{T}$ $\pi_{d}$ particles satisfy the following conditions: $0.2~\text{cm}<R_{xy}<60~\text{cm}$, $55~\text{cm}\leq z\leq 140~\text{cm}$, and $17^{\circ}\leq\theta_{\text{lab}}\leq 150^{\circ}$. Additionally, their daughter muons must satisfy $p_{T}>0.05~\text{GeV}$, and the reconstructed invariant mass of the muon pair must fulfill $m_{\mu^{+}\mu^{-}}<0.48~\text{GeV}$ or $m_{\mu^{+}\mu^{-}}>0.52~\text{GeV}$. In this analysis, we require at least two muon pairs with significantly displaced vertices, which effectively suppresses backgrounds from punch-through pions. For decay vertices with $R_{xy}>0.2~\text{cm}$, the significant displacement from the IP eliminates photon conversion backgrounds. Furthermore, backgrounds originating from detector structures can be effectively rejected for nonelectron final states. All signal events are required to satisfy at least one of the triggers listed in Table 2 and the selection criteria outlined in Table 3. Under these comprehensive selection criteria, the signal region can be treated as effectively background-free Duerr:2020muu ; Bernreuther:2022jlj .
@@
Table 4 summarizes the efficiencies of three trigger types and after applying event selection criteria for three BPs. Overall, the efficiencies of all three triggers exhibit significant variations with $m_{\pi_{d}}$, and the relative ordering of efficiencies among the BPs changes considerably after the same selection is applied. Specifically, the efficiency of the two tracks trigger decreases as $m_{\pi_{d}}$ increases, dropping from $75.82\%$ for BP1 to $34.16\%$ for BP3. This behavior can be understood from the distribution of $\theta_{\text{lab}}$. For $m_{\pi_{d}}=1~\text{GeV}$, the tracks from the leading and subleading $p_{T}$ $\pi_{d}$’s are more dispersed in $\theta_{\text{lab}}$ distribution, with many falling outside the trigger acceptance region of $38^{\circ}\leq\theta_{\text{lab}}\leq 127^{\circ}$. In this case, the two selected tracks are very likely to originate from the same dark pion, and their opening angle $\Delta\phi$ rarely exceeds $\pi/2$, leading to the lowest efficiency. In contrast, for $m_{\pi_{d}}=0.32$ and $0.7~\text{GeV}$, the $\theta_{\text{lab}}$ distribution is more concentrated, and the higher $\pi_{d}$ multiplicity results in higher efficiency. The one muon trigger maintains high efficiency across all three BPs, peaking at $89.22\%$ for BP2. The efficiency of the displaced vertex trigger decreases with increasing $m_{\pi_{d}}$. This condition strongly depends on the decay length of $\pi_{d}$. The proper decay lengths for the three BPs are $c\tau_{0}=116.40~\text{mm}$ (BP1), $8.80~\text{mm}$ (BP2), and $2.94~\text{mm}$ (BP3), respectively. The proper decay length of BP1 is an order of magnitude larger than that of BP2, after multiplying by $\beta\gamma$ in the lab frame, it often exceeds the upper limit of the trigger condition. However, since the proper decay length follows an exponential distribution in the Monte Carlo simulation, the high multiplicity of BP1 enhances the probability that at least one displaced vertex satisfies the trigger condition. In contrast, for $m_{\pi_{d}}=1~\text{GeV}$, the $R_{xy}$ distribution of the two leading $p_{T}$ $\pi_{d}$’s is concentrated in a very small region, resulting in the fewest events satisfying the trigger condition. After applying the cuts in Table 3, the order of efficiencies in these three cases may change. From the $R_{xy}$ distribution of $\pi_{d}$’s, for $m_{\pi_{d}}=0.32~\text{GeV}$, many events have decay lengths for both $\pi_{d}$’s exceeding 60 cm, leading to a significant reduction in efficiency. In comparison, for $m_{\pi_{d}}=0.7~\text{GeV}$, most events fall within the required range, resulting in the highest efficiency.
@@
We then analyze relevant signal kinematic distributions for $\pi_{d}$’s and muons at the Belle II in Figs. 7 and 8. First, as the $m_{\pi_{d}}$ increases from BP1 to BP3, the $N_{\pi_{d}}$ distribution (top left) decreases from approximately 6 to 2. Regarding the decay length, $R_{xy}$ and $R_{z}$ represent the transverse and longitudinal decay lengths of the dark pions in the lab frame, respectively. According to Eq. (4), under fixed parameters and with $f_{\pi_{d}}=m_{\pi_{d}}$, the proper decay length satisfies $c\tau_{0}\propto 1/m_{\pi_{d}}^{3}$. Therefore, lighter $\pi_{d}$’s have longer decay lengths, and their $R_{xy}$ (top right) and $R_{z}$ (bottom left) distributions exhibit larger values and longer tails. Finally, the $\theta_{\rm lab}(\text{track})$ distribution (bottom right) is significantly dispersed for the BP1, while the other two BPs exhibit relatively concentrated distributions. By comparing $E$ and $p_{T}$ of daughter particles, it can be seen that for BP1 and BP2, the leading and subleading $p_{T}$ $\pi_{d}$’s have similar $p_{z}$ values, leading to small differences in their $\Delta\theta_{\rm lab}$ distributions. In contrast, for the BP3, although its $p_{T}$ distribution is similar, the leading $\pi_{d}$ carries higher energy, resulting in a larger $p_{z}$ difference. The greater difference in boost further leads to a significant separation in their $\theta_{\rm lab}$ distributions.
@@
As shown in Fig. 8, the $p_{T}(\mu)$ distribution (top left) is similar to that in the BaBar analysis. In the $m_{\mu^{+}\mu^{-}}$ distribution (top right), the reconstructed mass peaks clearly appear near the respective $m_{\pi_{d}}$, $0.32$ GeV, $0.7$ GeV, and $1$ GeV for BP1, BP2, and BP3, respectively, but exhibit different levels of broadening. The width of these distributions is influenced by momentum resolutions of muons Adachi:2018qme ; Kang:2021oes ,
@@
the width of the $m_{\mu^{+}\mu^{-}}$ distribution increases from BP1 to BP3 as the momentum resolution grows with the $p_{T}(\mu)$. The $E_{\mu^{+}\mu^{-}}$ distribution (bottom left) shows a clear $m_{\pi_{d}}$ and $p_{T}(\mu)$ dependence-muon pairs from the decay of heavier and higher $p_{T}$ $\pi_{d}$’s carry more energy. Finally, in the $\Delta R_{\mu^{+}\mu^{-}}$ distribution (bottom right), the average angular separation between the muons is smaller for higher $\pi_{d}$ multiplicity, while in BP3, which has fewer $\pi_{d}$’s, the $\Delta R$ distribution is generally larger.
@@
Since the coupling $\kappa$ and mediator mass $m_{X}$ affect both the $e^{+}e^{-}\to\overline{q_{d}}q_{d}$ production cross section and $\pi_{d}$ lifetime, we map exclusion limits in the parameter space of $(m_{X},\kappa)$ plane. In particular, $\sigma(e^{+}e^{-}\to\overline{q_{d}}q_{d})$ and $\Gamma_{\pi_{d}}$ both scale as $(\kappa/m_{X})^{4}$. Applying the trigger conditions and event selection criteria (Tables 2 and 3), we obtain the signal efficiency $\epsilon_{S}$. Here an event is considered to pass the trigger selection if at least one of the three trigger conditions is satisfied. The expected signal yield is $N_{S}=\mathcal{L}\times\sigma(e^{+}e^{-}\to\overline{q_{d}}q_{d})\times\epsilon_{S}$, assuming optimistic integrated luminosity $\mathcal{L}=50~\text{ab}^{-1}$ for the Belle II experiment. With a background-free assumption, $N_{S}=2.3$ events establish the exclusion contours of the $90\%$ confidence level (C.L.) which correspond to the lower boundaries with purple solid lines in Fig. 6.
@@
As expected, small $\kappa/m_{X}$ yields a tiny cross section and excessive $\pi_{d}$ decay length, reducing $N_{S}$ below detection ability of the Belle II experiment. Conversely, large $\kappa/m_{X}$ produces prompt $\pi_{d}$ decays without the signature of displaced vertices. In this situation, the efficiency drops sharply, necessitating enormous Monte Carlo samples. Therefore, we conservatively set $N_{S}=10$ to estimate upper boundaries with purple dashed lines in Fig. 6333Even with this relaxed requirement, more than $10^{8}$ simulated events are required for this estimation.. We can observe that the projected Belle II constraints also indicate inverse proportionality between $m_{X}$ and $\kappa$ for these three BPs. Notably, blank regions in the upper-left corner of $(m_{X},\kappa)$ plane could be covered by searching for multimuons from the prompt decay of dark pions. However, detailed signal-background analysis is beyond the scope of this work.